# Relaxation time approximation revisited and pole/cut structure in retarded correlators

### v1

@@ III.2 Energy-dependent relaxation time approximation ($\alpha>0$)
As noted in the convenient parameterization Eq.(11), the relaxation time, the only parameter reflects the microscopic dynamic details within the transport model, is allowed to exhibit energy dependence. It has long been believed that incorporating various energy dependence can reveal characteristic features of bottom-up thermalization and uncover hidden aspects of the full kinetic description [15]. Furthermore, the linearized kinetic equation within the energy-dependent relaxation time approximation has been successfully applied to jet physics [29, 30], and hydrodynamic simulations [27, 28]. Does this imply that the model or approximation is well-justified? However, caution is warranted: an approximated model may capture some aspects of the underlying physics relevant to specific issues but may fail to do so for others. Nevertheless, elucidating how an approximation is derived from a complete theory is always highly significant. In this subsection, we show mathematically how this model can be developed from the linearized Boltzmann in a similar manner to what we have done in the former subsection, and how an inconsistence appears hindering the sound justification of the model.
@@ III.3 Energy-dependent relaxation time approximation ($\alpha<0$)
The lack of a well-defined relaxation time is due to the lack of a gap: As illustrated in Section III.2, a bounded energy-dependent relaxation time cannot be identified, but we can take a lesson from Eq.(11). It can be proved that such a gap can exist if the following three conditions are met:
@@
However, we contend that $\tau_{R}=\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$ equates to the scenario where $\alpha=0$, i.e., the AW RTA, and does not introduce a new class of RTAs. Suppose an observer perturbs the system in equilibrium with a small-amplitude spatially homogeneous disturbance at $t=0$, then he/she is observing when the disturbance dissipates to none. For the system described by Eq.(24), this relaxation scale is dictated by the mode with the longest lifetime, and reads $\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$. Hence at $t=\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$, the observer announces that the system returns to its equilibrium state. If the same experiment is conducted for the system Eq.(15), the observer should get the same result, as Eq.(24) is an equivalent transformation of Eq.(15). For future convenience in expression, this argument will be also referred to as the ”observer argument”. In other words, if condition 1 is satisfied, $\mathcal{L}_{0}$ also possesses nonzero eigenvalues separated from the origin, and these two relaxation scales should be identical, $\frac{1}{\gamma_{6}}\simeq\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$. Several comments on this relation are followed in order:
  • We identify a gapped scale $\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$, which bounds the relaxation timescales for perturbation decay. This means a RTA description emerges, with the relaxation time as given in Eq.(11) but with $E_{p}\rightarrow m$, which reminds me of an inverse but interesting replacement of $m\rightarrow E_{p}$ in the initial construction for AW RTA [21]. But ruled by the above physical arguments, this deformed energy-dependent RTA (precisely mass-dependent RTA) is still a redundant description. In this sense, we conclude that there is no energy-dependent RTA for $\mathcal{L}_{\alpha<0}$, but just a trivial mass-dependent RTA.
@@ III.4 Hard interaction VS soft interaction
where $\nu(p)$ represents the collision frequency and we introduce a compact operator $K$. Here we maintain the same notations as in the previous context to avoid complicating the overall notation system although Eq.(26) looks quite different. Furthermore, it is immediately observable that Eq.(27) shares the same form as Eq.(10). As a result, they should also share the same formal solution $\chi(t)=e^{-\mathcal{L}_{0}t}\chi(0)$ when the system is homogeneously perturbed. In both cases, whether considering Eq.(27) or Eq.(10), $\chi$ should be regarded as the perturbation above the same equilibrium state. Given that both Eqs.(27) and (10) are derived from the same Boltzmann equation through linearization, we would not expect the eigenspectrum for $\mathcal{L}_{0}$ in Eq.(27) to differ from that in Eq. (10). If this were the case, the decay behavior of the perturbation would exhibit significant differences, signifying the physical inconsistency. This is another successful application of the observer argument from the previous subsection. Therefore, we argue that the eigenspectrum of $\mathcal{L}_{0}$ should be identical in Eqs.(27) and (10), and therefore we maintain the same notations. Then let’s come to the core of this subsection:
@@
Proofs of these theorems can be found in [32]. Analyzing the collision frequency $\nu(p)$ enables us to distinguish between relativistic soft and hard interactions. Mathematical analysis indicates that relativistic interactions tend to be softer than their nonrelativistic counterparts: there is a large class of functions $\sigma(g,\theta)$, which corresponds to scattering cross-sections for relativistic soft, but non-relativistic hard interactions [31, 32]. It can be also proved that for soft interactions $-\mathcal{L}_{0}$ is a bounded operator with the eigenspectrum $[-\nu_{max},0]$, whereas for hard interactions, $-\mathcal{L}_{0}$ is unbounded, featuring an eigenspectrum of $[-\infty,-\nu_{min}]$, where $\nu_{max}$ and $\nu_{min}$ can be estimated by the extrema of $\nu(p)$. Given the eigenspectrum properties presented here, we conclude that the AW RTA is only well-justified in the case of hard interactions. As most relativistic interactions are soft according to Theorem 2, the AW RTA lacks a solid foundation in the majority of cases. For instance, in the extensively studied scalar $\phi^{4}$ theory, where $\sigma(g,\theta)\sim\frac{1}{s}=\frac{1}{g^{2}+4m^{2}}<\frac{1}{g^{2}}$, the interaction falls within the soft interaction category (this is confirmed in numerical calculation [6]). This softness results in a branch-cut structure that extends across the entire negative imaginary axis in the retarded correlators [2, 33]. In the next section, the relationship between Q1 and Q2 will be clarified based upon the insights gained from our previous discussions.
@@ III.5 Looking for RTA and eigenspectrum properties of $\mathcal{L}_{\alpha}$
The observer argument can not only rule out redundant RTAs, but also constrain certain eigenspectrum properties of $\mathcal{L}_{\alpha\neq 0}$, given the softness of the interactions, as will be detailed below. One may wonder if it is necessary to pay attention to the properties of $\mathcal{L}_{\alpha\neq 0}$ eigenspectrum, especially given that $\mathcal{L}_{0}$ is more naturally associated with the analytical properties of retarded correlators exhibited in the next section. In principle, the selection of $\mathcal{L}_{\alpha}$ is not fixed; one can select any suitable option. On one hand, it is more suitable to start from $\mathcal{L}_{\alpha\neq 0}$ sometimes. For instance, a recent calculation on the eigenspectrum of $\mathcal{L}_{1}$ within the massless scalar $\phi^{4}$ theory was given in an analytical way [5], while it is more challenging to perform such calculations for $\mathcal{L}_{0}$ (this may also be the case for other interactions). On the other hand, $\mathcal{L}_{1}$ naturally appears on the right-hand side of the linearized Boltzmann equation in Eq.(3) if no further transformations are made [24]. To sum up, we think that discussing the properties of $\mathcal{L}_{\alpha\neq 0}$ eigenspectrum is of profound importance.
Then let’s elaborate on how the analysis is implemented. We opt to present this result via a case-by-case discussion, and remind that ”the eigenspectrum ($\gamma_{n}$ or $\gamma^{\prime}_{n}$ for eigenvalues) is gapped or not ” and ”the decay/relaxation rates, the inverse of the perturbation’s relaxation timescales $\frac{\gamma_{n}}{E_{p}^{\alpha}}$, are gapped or not ” are entirely distinct concepts.
@@
    • $m\neq 0$: The eigenspectrum of $\mathcal{L}_{\alpha>0}$ is still gapless for the same reason. However, the relaxation timescales for $\mathcal{L}_{\alpha<0}$ are bounded by $\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$ provided that the eigenspectrum of $\mathcal{L}_{\alpha<0}$ is well-separated from zero. In order to match the gapless nature of $\mathcal{L}_{0}$, the eigenspectrum of $\mathcal{L}_{\alpha<0}$ must be also gapless, unless the observer argument is violated.
@@
    • $m=0$: As mentioned above, the presence of $E_{p}^{\alpha}(\alpha\neq 0)$ potentially rendering the relaxation rates gapless, contradicting the gapped rates for $\mathcal{L}_{0}$. To restore the physical observer argument, we propose the following conditions that could resolve the potential conflicts: the eigenspectrum of $\mathcal{L}_{\alpha\neq 0}$ is gapped; the eigenfunctions are localized within a finite $p$ region $p\leq p_{max}$ for $\mathcal{L}_{\alpha>0}$; the eigenfunctions are bounded in a region $p\geq p_{min}>0$ for $\mathcal{L}_{\alpha<0}$. As can be seen clearly, the argument imposes rather strong constraints if the guessed conditions are true, and furthermore, $\frac{1}{\gamma_{6}}\simeq\frac{p_{max}^{\alpha}}{\gamma^{\prime}_{6}}$ for $\alpha>0$ and $\frac{1}{\gamma_{6}}\simeq\frac{p_{min}^{\alpha}}{\gamma^{\prime}_{6}}$ for $\alpha<0$. We propose a testing method to verify or falsify the above guesses: numerically solving the eigenspectrum of $\mathcal{L}_{\alpha\neq 0}$ for hard interactions [6], for instance, four fermions interaction in the electroweak sector.
@@ IV Pole or cut  — the non-analytical structures in retarded correlators
Two-point retarded correlation functions are crucial and insightful as they encapsulate rich information about the transport properties of many-body systems. For instance, their analytical structure can reflect the characteristic properties of how thermal equilibrium is reached. In addition, their non-analytical structures — poles or cuts in Fourier space, governs the evolution behavior of the system: poles describe collective excitations evolving towards equilibrium corresponding to the longtime hydrodynamic modes, while the presence of cuts or non-hydrodynamic modes is closely linked to the emergence of hydrodynamic behavior and the applicability of hydrodynamics. Thus, research into the analytical properties of retarded correlators is profound, which is initially explored in the weakly-coupled kinetic theory by Romatschke [1]. This work was motivated by attempts to comprehend calculations in strongly-coupled gauge theories with holographic duals, like $\mathcal{N}=4$ SYM theory at strong coupling with many colors. Romatschke’s findings highlight two key features in the analytical properties of the retarded correlators: the cuts are gapped corresponding to non-hydrodynamic modes, below which hydrodynamic poles dominate; the hydrodynamic poles cease to exist for some critical value of the wavenumber reminiscent of the phenomenon of onset transitions, which are successfully reproduced within the mutilated RTA model detailed in [26]. As a supplement, we note that the universal behavior of onset transitions has been reported for a long time in the context of the nonrelativistic linearized Boltzmann equation using the mutilated model [20]. However, the AW RTA adopted in [1] is less well-founded than its non-relativistic counterpart, given that interactions tend to be softer in relativistic scenarios: the universal onset transitions in nonrelativistic systems may be rarely observed in relativistic systems due to varying degrees of interaction softness.
@@
Later on, Kurkela and Wiedemann reexamined the behavior of the retarded correlators, beginning with the parametrized energy-dependent RTA as described in Eq. (11), with a particular focus on the case when $\alpha=1$ [15]. Their conclusions, however, stand in contrast to those in [1]: there is no sharp onset of hydrodynamic behavior; the structure of cuts turns into the entire strip $\text{Im}\,\omega<0,-k\leq\text{Re}\,\omega\leq k$ from a gapped line in [1]. They also state that the appearance of poles in the first (physical) Riemann sheet of retarded correlation functions is a matter of choosing a particular analytical continuation and thus cannot be related unambiguously to the onset of fluid dynamic behavior. However, as demonstrated in III.2, the model corresponding to Eq.(11) with $\alpha=1$ exhibits inconsistencies: it can only be seen as an incomplete truncation to its ultraviolet (UV) completion — the linearized Boltzmann equation, because an infinite number of slow modes are excluded for lack of the only slowest mode. Nevertheless, the analysis given in [15] is still illuminating, and qualitatively predicts the cut structure within the stress-stress retarded correlators, namely, a cut along the negative imaginary frequency axis when $k=0$ [2]. We note that, although the model is incomplete, the retained part may still contain sufficient information to reflect the non-analytical structures within the retarded correlators, depending on the nature of the interactions. We provide evidence for this assertion based on recent calculations in [4], where the authors demonstrate that the shear-stress linear response function exhibits a branch-cut singularity extending across the entire negative imaginary semi-axis. This is achieved using an analytical result for the eigensystem of the linearized collision term in the massless scalar $\phi^{4}$ theory with leading-order interaction.
@@
As observed, these two studies introduced above correspond to the models discussed in III.1 and III.2, thereby clarifying the relationship between Q1 and Q2. In this place, we want to give a general statement on the subject of pole/cut structure in the (stress-stress) retarded correlators. Before proceeding, let’s elaborate on how to derive the non-analytical structure if the interactions are soft. In this case, the eigenvalue spectrum of $\mathcal{L}_{0}$ is gapless and the typical stress-stress retarded correlation function is given by
@@
In the limit of vanishing mass and nonzero $k$, we replace $\omega$ by $\omega-k\cos\theta$ where $\theta$ is the angle between $p$ and $k$. Then we can cast $G_{R}(\omega)$ into
@@
where Eq.(37) is made use of in advance $\lim\limits_{m\rightarrow 0}\nu_{max}(m)\rightarrow\infty$. Ultimately, we successfully derive the branch-cut structure for nonzero $k$  [^2] While [15] provides the correct structure for nonzero $k$, it is based on the choice of $\mathcal{L}_{1}$, or more precisely, an incomplete truncation of $\mathcal{L}_{1}$, as shown in III.2. For completeness and consistency, a thorough study extending [4] to spatially inhomogeneous perturbations is warranted, which should also give Eq.(34)..
This result can also be intuitively understood by comparing it with Romatschke’s RTA calculation. Recalling that the RTA used in [1] truncates the full operator into a single value, we should retain a series of eigenvalues near the origin to ”model” the soft interactions. Note that a single eigenvalue or a single relaxation time results in a branch-cut line
$\text{Im }\omega=-\frac{1}{\tau_{R}},\quad-k\leq\text{Re}\,\omega\leq k.$ (35)
When accounting for additional eigenvalues, more branch-cut lines emerge. If these eigenvalues form a continuous spectrum, then the following strip
$-\infty<\text{Im }\omega<0,\quad-k\leq\text{Re}\,\omega\leq k$ (36)
can be recovered. In the homogeneous limit $k\to 0$, such a strip shrinks to a branch-cut line extending across the negative imaginary axis.
Based on the above statement, determining the pole/cut structure in the retarded correlation functions equates to solving for the eigenspectrum of $\mathcal{L}_{0}$, thereby interlinking Q1 and Q2. However, a contradictory question arises: both [2] and [6] demonstrate that the cut extends across the entire negative imaginary frequency axis, contrasting with the bounded region $[-i\,\nu_{max},0]$ inferred from mathematical analysis. The discrepancy vanishes because, in the context of a massless theory with soft interactions, the collision frequency becomes unbounded [32]
$\nu(p)\longrightarrow\infty,\quad\text{when}\quad p\rightarrow 0$ (37)
where $C$ is an irrelevant constant. Therefore, $\nu_{max}=\infty$ and these results are consistent with each other. We give a short summary of this section by reconsidering the interplay of nonzero particle mass $m$ and wavenumber $k$, and our statement can be summarized as follows:
@@
    • $m=0$: Romatschke’s analysis applies: the retarded correlators (including stress-stress correlator) exhibit gapped branch-cut lines Eq.(35) associated with nonhydrodynamic modes. Below this gap, hydrodynamic modes represent the sole low-energy degrees of freedom, indicating that hydrodynamics is well-defined in such scenarios.
    • $m\neq 0,\,k=0$: The branch-cut line turns into a branch point $\text{Im }\omega=-\frac{1}{\tau_{R}},\,\text{Re}\,\omega=0$.
    • $m\neq 0,\,k\neq 0$: The non-analytical structures of the retarded correlators become complicated due to the interplay between the nonzero particle mass $m$ and wavenumber $k$. Upon inspecting the derivation detailed in [1], we find that if the particles are massive, the term proportional to $\bm{v}\cdot\nabla\sim\frac{pk\cos\theta}{\sqrt{p^{2}+m^{2}}}$ depending on $p$. This could introduce additional complex structures into correlators after performing momentum integral, beyond the poles or gapped cuts predicted by Romatschke. In the limit of vanishing mass or $k$, the conclusions align with the above two cases. For now, we have no definite results for $m\neq 0,\,k\neq 0$, but we leave the analytical calculation to an ongoing task. Note as an aside, a numerical calculation [6] regarding four fermions interaction in the electroweak sector or massive RTA would be helpful to elucidate this issue.
@@
      which is exactly the result reported in [2].
@@
    • $m=0,\,k\neq 0$: The branch-cut of the stress-stress correlator has been extended to the case of nonzero $k$ in Eq.(34)
      $-\infty<\text{Im }\omega<0,\quad-k\leq\text{Re}\,\omega\leq k.$ (40)
    • $m\neq 0,\,k\neq 0$: The complication arising from $\frac{pk\cos\theta}{\sqrt{p^{2}+m^{2}}}$ also exists in this scenario, and the issue should be trickier. If analytical treatment is not a strict requirement, the numerical technique in [6] would be helpful. In this scenario, we can examine the simplest interaction case, massive scalar $\phi^{4}$ theory with leading-order interaction. Our efforts in this direction are in progress.

### v2

@@ III.2 Energy-dependent relaxation time approximation ($\alpha>0$)
As noted in the convenient parameterization Eq.(8), the relaxation time, the only parameter reflecting the microscopic dynamic details within the RTA model, is allowed to exhibit energy dependence. It has long been believed that incorporating various energy dependence can reveal characteristic features of bottom-up thermalization and uncover hidden aspects of the full kinetic description [10]. Furthermore, the linearized kinetic equation within the energy-dependent relaxation time approximation has been successfully applied to jet physics [28, 29], hydrodynamic simulations [25, 26], and Weyl semimetals [30]. Does this imply that the model or approximation is well-justified? Caution should be taken: an approximated model may capture some aspects of the underlying physics relevant to specific issues but may fail to do so for others. Nevertheless, elucidating how an approximation is derived from a complete theory is always highly significant. In this subsection, we show mathematically how this model can be developed from the linearized Boltzmann in a similar manner to what we have presented in the former subsection, and how an inconsistence appears hindering the sound justification of the model.
@@ III.3 Energy-dependent relaxation time approximation ($\alpha<0$)
The lack of a well-defined relaxation time is due to the lack of a gap: As illustrated in Section III.2, a bounded energy-dependent relaxation time cannot be identified therein, but we can take a lesson from Eq.(8). It can be proved that such a gap can exist if the following three conditions are met:
@@
However, we confirm that $\tau_{R}=\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$ equates to the scenario where $\alpha=0$, i.e., the AW RTA, and does not introduce a new class of RTAs. Suppose an observer perturbs the system in equilibrium with a small-amplitude spatially homogeneous disturbance at $t=0$, then the observer is observing when the disturbance dissipates to vanishing. For the system described by Eq.(21), this relaxation scale is dictated by the mode with the longest lifetime, and reads $\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$. Hence at $t=\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$, the observer announces that the system returns to its equilibrium state. If the same experiment is conducted for the system Eq.(12), the observer should get the same result, as Eq.(21) is an equivalent transformation of Eq.(12). For future convenience in expression, this argument will be also referred to as the “observer argument”. In other words, if condition 1 is satisfied, $\mathcal{L}_{0}$ also possesses nonzero eigenvalues separated from the origin, and these two relaxation scales should be identical, $\frac{1}{\gamma_{6}}\simeq\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$. Several comments on this relation are followed in order:
  • We identify a gapped scale $\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$, which bounds the relaxation timescales for perturbation decay. This means a RTA description emerges, with the relaxation time as given in Eq.(8) but with $E_{p}\rightarrow m$, which reminds me of an inverse but interesting replacement of $m\rightarrow E_{p}$ in the initial construction for AW RTA [20]. But ruled by the above physical arguments, this deformed energy-dependent RTA (precisely mass-dependent RTA) is still a redundant description. In this sense, we conclude that there is no energy-dependent RTA for $\mathcal{L}_{\alpha<0}$.
@@ III.4 Hard interaction VS soft interaction
where $\nu(p)$ represents the collision frequency and we introduce a compact operator $K$. Here we maintain the same notations as in the previous context to avoid complicating the overall notation system, although the expansion Eq.(23) looks quite different. Furthermore, it is immediately observable that Eq.(24) shares the same form as Eq.(7). As a result, they should also share the same formal solution $\chi(t)=e^{-\mathcal{L}_{0}t}\chi(0)$ when the system is homogeneously perturbed. In both cases, whether considering Eq.(24) or Eq.(7), $\chi$ should be regarded as the perturbation above the same equilibrium state. Given that both Eqs.(24) and (7) are derived from the same Boltzmann equation through linearization, we would not expect the eigenspectrum for $\mathcal{L}_{0}$ in Eq.(24) to differ from that in Eq. (7). If this were the case, the decay behavior of the perturbation would exhibit significant differences, signifying the physical inconsistency. This is another successful application of the observer argument from the previous subsection. Therefore, we argue that the eigenspectrum of $\mathcal{L}_{0}$ should be identical in Eqs.(24) and (7), and this is why we maintain the same notations. Then let’s come to the core of this subsection:
@@
Proofs of these theorems can be found in [32]. Analyzing the collision frequency $\nu(p)$ enables us to distinguish between relativistic soft and hard interactions. Mathematical analysis indicates that relativistic interactions tend to be softer than their nonrelativistic counterparts: there is a large class of functions $\sigma(g,\Theta)$, which corresponds to scattering cross-sections for relativistic soft, but non-relativistic hard interactions [31, 32]. It can be also proved that for soft interactions $-\mathcal{L}_{0}$ is a bounded operator with the eigenspectrum $[-\nu_{max},0]$, whereas for hard interactions, $-\mathcal{L}_{0}$ is unbounded, featuring an eigenspectrum of $[-\infty,-\nu_{min}]$, where $\nu_{max}$ and $\nu_{min}$ can be estimated by the extrema of $\nu(p)$. Given the eigenspectrum properties presented here, we conclude that the AW RTA is only well-justified in the case of hard interactions. As most relativistic interactions are soft according to Theorem 2, the AW RTA lacks a solid foundation in the majority of cases. For instance, in the extensively studied scalar $\phi^{4}$ theory, where $\sigma(g,\Theta)\sim\frac{1}{s}=\frac{1}{g^{2}+4m^{2}}<\frac{2}{g^{2}}$, the interaction falls within the soft interaction category (this is confirmed in numerical calculation [13]). This softness results in a branch-cut structure that extends across the entire negative imaginary axis in the retarded correlators [12, 33]. In the next section, the relationship between Q1 and Q2 will be clarified based upon the insights gained from our previous discussions.
@@ III.5 Looking for RTA and eigenspectrum properties of $\mathcal{L}_{\alpha}$
The observer argument can not only rule out redundant RTAs, but also constrain certain eigenspectrum properties of $\mathcal{L}_{\alpha\neq 0}$, given the softness of the interactions, as will be detailed below. One may wonder if it is necessary to pay attention to the properties of $\mathcal{L}_{\alpha\neq 0}$ eigenspectrum, especially given that $\mathcal{L}_{0}$ is more naturally associated with the analytical properties of retarded correlators, which will be exhibited in the next section. In principle, the selection of $\mathcal{L}_{\alpha}$ is not fixed; one can select any suitable option. It is more suitable to start from $\mathcal{L}_{\alpha\neq 0}$ sometimes. For instance, a recent calculation on the eigenspectrum of $\mathcal{L}_{1}$ within the massless scalar $\phi^{4}$ theory was given in an analytical way [18], while it is more challenging to perform such calculations for $\mathcal{L}_{0}$ (this may also be the case for other interactions). Therefore, we think that discussing the properties of $\mathcal{L}_{\alpha\neq 0}$ eigenspectrum is not trivial at all.
Then let’s elaborate on how the constraints can be imposed. We opt to present the results via a case-by-case discussion, and remind that “the eigenspectrum ($\gamma_{n}$ or $\gamma^{\prime}_{n}$ for eigenvalues) is gapped or not” and “the decay/relaxation rates, the inverse of the perturbation’s relaxation timescales $\frac{\gamma_{n}}{E_{p}^{\alpha}}$, are gapped or not” are entirely distinct concepts.
@@
    • $m\neq 0$: The eigenspectrum of $\mathcal{L}_{\alpha>0}$ is still gapless for the same reason. For $\alpha<0$, the relaxation timescales are bounded by $\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$ provided that the eigenspectrum of $\mathcal{L}_{\alpha<0}$ is gapped. However, to match the gapless nature of $\mathcal{L}_{0}$, the eigenspectrum of $\mathcal{L}_{\alpha<0}$ must be also gapless.
@@
    • $m=0$: As mentioned above, the presence of $E_{p}^{\alpha}(\alpha\neq 0)$ potentially rendering the relaxation rates gapless, contradicting the gapped relaxation rates for $\mathcal{L}_{0}$. To restore the physical observer argument, we propose the following conditions that could resolve the potential conflicts: the eigenspectrum of $\mathcal{L}_{\alpha\neq 0}$ is gapped; the eigenfunctions are localized within a finite $p$ region $p\leq p_{max}$ for $\mathcal{L}_{\alpha>0}$; the eigenfunctions are bounded in a region $p\geq p_{min}>0$ for $\mathcal{L}_{\alpha<0}$. As can be seen clearly, the argument imposes rather strong constraints, and furthermore, $\frac{1}{\gamma_{6}}\simeq\frac{p_{max}^{\alpha}}{\gamma^{\prime}_{6}}$ for $\alpha>0$ and $\frac{1}{\gamma_{6}}\simeq\frac{p_{min}^{\alpha}}{\gamma^{\prime}_{6}}$ for $\alpha<0$. We propose a testing method to verify or falsify the above assumed constraints: numerically solving the eigenspectrum of $\mathcal{L}_{\alpha\neq 0}$ for hard interactions [13], for instance, four fermions interaction in the low-energy electroweak sector.
@@ IV Pole or cut  — the non-analytical structures in retarded correlators
Two-point retarded correlation functions are crucial and insightful as they encapsulate rich information about the transport properties of many-body systems. Their analytical structures can reflect the characteristic properties of how thermal equilibrium is reached. For instance, the non-analytical structures — poles or cuts in Fourier space, governs the evolution behavior of the system: poles describe collective excitations evolving towards equilibrium corresponding to hydrodynamic modes, while the presence of cuts or non-hydrodynamic modes is closely linked to the emergence of hydrodynamic behavior and the applicability of hydrodynamics. Thus, the research into the analytical properties of retarded correlators is profound, which is initially explored by Romatschke in the weakly-coupled kinetic theory [7]. Romatschke’s findings highlight two key features in the analytical properties of the retarded correlators: the cuts are gapped corresponding to non-hydrodynamic modes, below which hydrodynamic poles dominate as long-lived degrees of freedom; the hydrodynamic poles cease to exist for some critical value of the wavenumber reminiscent of the phenomenon of onset transitions, which are successfully reproduced within the mutilated RTA model detailed in [34]. As a supplement, we note that the universal behavior of onset transitions has been reported for a long time in the context of the nonrelativistic kinetic theory using the mutilated model [19]. However, the AW RTA adopted in [7] is less well-founded than its non-relativistic counterpart, given that interactions tend to be softer in relativistic scenarios: the universal onset transitions in nonrelativistic systems may be rarely observed in relativistic systems due to varying degrees of interaction softness.
@@
Later on, Kurkela and Wiedemann reexamined the behavior of the retarded correlators, beginning with the parametrized energy-dependent RTA as described in Eq. (8), with a particular focus on the case when $\alpha=1$ [10]. Their conclusions, however, stand in contrast to those in [7]: there is no sharp onset of hydrodynamic behavior; the structure of cuts turns into the entire strip $\text{Im}\,\omega<0,-k\leq\text{Re}\,\omega\leq k$ from a gapped line given in [7]. They also state that the appearance of poles in the first (physical) Riemann sheet of retarded correlation functions is a matter of choosing a particular analytical continuation and thus cannot be related unambiguously to the onset of fluid dynamic behavior. However, as demonstrated in III.2, their model corresponding to Eq.(8) with $\alpha=1$ exhibits inconsistencies: it can only be seen as an incomplete truncation to its ultraviolet (UV) completion — the linearized Boltzmann equation, because an infinite number of slow modes are excluded. Although the analysis given in [10] is still illuminating, we choose to work with RTA only in the hard interaction case for theoretical consistence.
@@
As observed, the above two studies introduced above correspond to the models discussed in III.1 and III.2, thereby clarifying the relationship between Q1 and Q2. In this place, we want to give a general statement on the subject of pole/cut structure in the (stress-stress) retarded correlators. Before proceeding, let’s elaborate on how to derive the non-analytical structure if the interactions are soft. In this case, the eigenvalue spectrum of $\mathcal{L}_{0}$ is gapless and the typical stress-stress retarded correlation function is given by
@@
In the limit of vanishing mass and nonzero $k$, extracting the non-analytical structure becomes unmanageable. If $\cos\theta$, regarded as an operator in momentum-space functions, commute with $\mathcal{L}_{0}$, we can replace $\omega$ by $\omega-k\cos\theta$, where $\theta$ is the angle between spatial components of $p$ and $k$. Then we can cast $G_{R}(\omega)$ into
@@
where the reason for $\nu_{max}=0$ is postponed to Eq.(32). Ultimately, we have successfully reproduce the branch-cut structure for nonzero $k$ given in [10] 222 [10] is based on the choice of $\mathcal{L}_{1}$, or more precisely, an incomplete truncation of $\mathcal{L}_{1}$, as shown in III.2. For completeness and consistency, a thorough study extending [35] to spatially inhomogeneous perturbations is warranted.. This accordance is expected because $[\cos\theta,\mathcal{L}_{\text{RTA}}]=0$.
However, a reminder should be given that Eq.(31) relies heavily on the commutation approximation for $\cos\theta$ and $\mathcal{L}_{0}$, which breaks down when $\mathcal{L}_{0}$ is the complete linearized Boltzmann collision operator rather than the RTA. In other words, the structure in Eq.(31) should be modified. For situations involving hard interactions, the extent of modification should be minimal as the RTA is well-defined in that case.
Based on the above statement, determining the non-analytical structures in the retarded correlation functions equates to solving for the eigenspectrum of $\mathcal{L}_{0}$, thereby interlinking Q1 and Q2. However, a contradictory question arises: both [12] and [13] demonstrate that the cut extends across the entire negative imaginary frequency axis, contrasting with the bounded region $[-i\,\nu_{max},0]$ inferred from mathematical analysis. The discrepancy vanishes because, in the context of a massless theory with soft interactions, the collision frequency becomes unbounded [32]
$\nu(p)\longrightarrow\infty,\quad\text{when}\quad p\rightarrow 0.$ (32)
Therefore, $\nu_{max}=\infty$ and these results are consistent with each other.
For clarity, we give a short summary of this section by reconsidering the interplay of nonzero particle mass $m$ and wavenumber $k$, and our statement can be summarized as follows:
@@
    • $m=0$: Romatschke’s analysis applies: the retarded correlators (including stress-stress correlator) exhibit gapped branch-cut lines
      $\text{Im }\omega=-\frac{1}{\tau_{R}},\quad-k\leq\text{Re}\,\omega\leq k$ (33)
      associated with nonhydrodynamic modes. Below this gap, hydrodynamic modes represent the sole low-energy degrees of freedom, indicating that hydrodynamics is well-defined in such scenarios.
    • $m\neq 0,\,k=0$: The branch-cut lines turn into branch points $\text{Im }\omega=-\frac{1}{\tau_{R}},\,\text{Re}\,\omega=0$.
    • $m\neq 0,\,k\neq 0$: The non-analytical structures of the retarded correlators become complicated due to the interplay between the nonzero particle mass $m$ and wavenumber $k$. Upon inspecting the derivation detailed in [7], we find that if the particles are massive, the free-streaming is proportional to $\bm{v}\cdot\nabla\sim\frac{pk\cos\theta}{\sqrt{p^{2}+m^{2}}}$ depending on $p$. This could introduce additional complex structures into correlators after performing momentum integral, beyond the poles or gapped cuts predicted by Romatschke. For now, we cannot exclude the possibility of closing the window where hydrodynamic modes are the only long-lived degrees of freedom. In the limit of vanishing mass or $k$, the conclusions align with the above two cases.
@@
      which matches the conclusion given in [12]. Strictly speaking, the author focus on the discussions on the symmetrized 2-point function therein, but the retarded correlation function can be related to it through KMS relation. At weak coupling where the relevant frequencies will be suppressed by powers of the coupling, the result aligns with Eq.(34), which is further confirmed in a related study discussing $\mathcal{L}_{1}$ [35].
@@
    • $k\neq 0$: The complication arising from $\frac{pk\cos\theta}{\sqrt{p^{2}+m^{2}}}$ also exists in this scenario, and the issue should be trickier. Even if $m=0$, it is still difficult to reach a concise-form conclusion due to the non-commutativity of $\cos\theta$ and $\mathcal{L}_{0}$.
Although failing to reach a definite conclusion in the cases involved with nonzero $k$ or/and $m$, we plan to numerically solve the issue following techniques developed in [13], which is left to a future work. For instance, for soft interactions with nonzero $k$ or/and $m$, we can examine the simplest interaction case, scalar $\phi^{4}$ theory with leading-order interaction. As a representative theory within the category of soft interactions, this calculation is expected to reveal the property of non-analytical structures qualitatively. Our numerical efforts following this direction are in progress.